OpenStudy (anonymous):

What is the solution to the integral below:

OpenStudy (anonymous):

where is it?

OpenStudy (anonymous):

\[\int\limits_{1}^{x} (2/x^2)\]

OpenStudy (cybershadow):

see it wud be 2(x^-2) inside the integral

OpenStudy (cybershadow):

as 2 is the constant we can take it out and therefore it is (x^-2) inside the integral

OpenStudy (cybershadow):

that will make it 2 * ( x^(-2+1)/-2+1)

OpenStudy (cybershadow):

put in the limits and solve it :)

OpenStudy (anonymous):

Sorry I was gone for a second

OpenStudy (anonymous):

Oh I see. We'll get 2x^-1/-1 = -2x^-1

OpenStudy (anonymous):

+2

OpenStudy (cybershadow):

u forgot to insert the upper limit

OpenStudy (cybershadow):

it wud be 2(x^-1 -1)

OpenStudy (anonymous):

Oh okay.

OpenStudy (cybershadow):

yes and to write it more simpler ,u can write it as 2(1/x-1) = 2(1-x/x)